OpenStudy (anonymous):

Edward borrows $350 and pays 3.5% simple intrest each year. If he pays back the money in 1 year, what is the total amount he pays?

OpenStudy (anonymous):

@Directrix

OpenStudy (anonymous):

@Preetha

Directrix (directrix):

Interest = P*R*T where p is the principal, r is the interest rate, and t is the time Ed will owe 350 + .035*350 at the end of the year. He owes $ 362.65

OpenStudy (anonymous):

Interest =P*r*t where Pis the principal, r is the interest rate per term and t is the number of terms I=P*r*t P=$350 r=3.5*100=0.35p.a t=1 year I=350*0.35*1 =122.50 The amount of interest he would pay is $122.50. however since it is asking for the total amount that you are looking for there is another step that needs to be done. A=P+I where A is the total amount due at the end of the investment P is the principal the amount of money invested I is the simple interest earned A=P+I P=$350 I=$122.50 A= 350+122.50 A=$472.50 He would owe $472.50. I could be wrong about that last step it could be just the interest. I hope this helps you.

Directrix (directrix):

@renni3 3.5% = .035 as a decimal, not 0.35 . I think that is why our answers differ.

OpenStudy (anonymous):

oops I think you are right there. sorry

OpenStudy (anonymous):

Thank you @renni3 & @Directrix so much!!!!!